A is 4 rowed square matrix. |A|=4, adj(adj(A))=? a)2A b)4A c)8A d)16A
1
Expert's answer
2014-01-20T03:10:55-0500
Using Adjugate matrix properties we have
adj(adj(A))=det(A)^(n-2)A.
In our case n=4 and det(A)=4. Substitute these into previous formula
adj(adj(A))=4^(4-2)A=16A.
Therefore, d) is the right answer.